Please help me solve the below:
Bruce has $5.15 in dimes and quarters. He has 8 more quarters than dimes. How many quarters does he have ??
thank you !!!!!
(HS Freshman struggling in Algebra)

Answer by TutorDelphia(193)

I like doing this problems in terms of pennies rather than dollars.
So a dime is worth 10 and quarter is worth 25. That gives me this equation:
10d+25q=515
He has 8 more quarters than dimes can be written as:
d+8=q
This is already solved for q so I can sub in d+8 for q since they have equal values
That gives me:
10d+25(d+8)=515
Distribute the 25
10d+25d+200=515
Combine like terms and subtract 200 from both sides
35d=315
Divide by 35
d=9
Sub in 9 for d
9+8=q
q=17
If you need more help, just let me know.
Thanks,
Adam